Our hypothetical population contains the scores 4, 6, 7, and 9. Determine


the mean and variance of the sampling distribution of the sample mean,


given that samples contain two scores drawn from the population with


replacement?



1
Expert's answer
2022-03-08T05:20:02-0500

Let X be the random variable representing the possible values of the sample means

Given the  the scores 4, 6, 7, and 9, the possible samples and their means are:

Then,

Sample space Mean (X)

{4,4} (4+4)/2 = 4

{4,6} (4+6)/2 = 5

{6,4} (6+4)/2 = 5

{4,7} (4+7)/2 = 5.5

(7,4} (7+4)/2 = 5.5

{4,9} (4+9)/2 = 6.5

{9,4} (9+4)/2 = 6.5

{6,6) (6+6)/2 = 6

{6,7} (6+7)/2 = 6.5

{7,6} (7+6)/2 = 6.5

{6,9} (6+9)/2 = 7.5

{9,6} (9+6)/2 = 7.5

{7,9} (7+9)/2 = 8

{9,7) (9+7)/2 = 8

{7,7) (7+7)/2 = 7

{9,9} (9+9)/2 = 9


Therefore, the possible values for X with respective probabilities are:

X P(X)

4 1/16

5 2/16

5.5 2/16

6 1/16

6.5 4/16

7 1/16

7.5 2/16

8 2/16

9 1/16


Then, the mean of X = ΣX*P(X)

=4(1/16)+5(2/16)+5.5(2/16)+6(1/16)+6.5(4/16)+7(1/6)+7.5(2/16)+8(2/16)+9(1/16)

=7.23

Answer: Mean of X = 7.23


Variance of X

σX2 = Σ(x – μ)2⋅ P(x)

="\frac{1}{6}"(4-7.23)2 + "\frac{2}{16}"(5-7.23)2 + "\frac{2}{16}"(5.5-7.23)2"\frac{1}{16}"(6-7.23)2 + "\frac{4}{16}"(6.5-7.23)2 + "\frac{1}{16}"(7-7.23)2 +"\frac{2}{16}"(7.5-7.23)2"\frac{2}{16}"(8-7.23)2"\frac{1}{16}"(9-7.23)2


= 2.1579

Answer: Variance of X = 2.1579